OpenStudy (anonymous):

the decimal from 1.46 becomes expressed as a percent

OpenStudy (mokeira):

1.46 is the same as \[\frac{ 146 }{ 100 }\] simply multiply by 100 to get percentage \[\frac{ 146 }{100 } \times 100\] 146%

OpenStudy (mokeira):

@shatiaray22 have you understood?